Question 28 0 out of 2 po n a case-control study of osteoporosis, 150 women were enrolled as cases and an equal number as controls. Among the cases 50 reported low calcium intake, while only 20 of the controls reported low calcium intake. Calculate the odds ratio in the association between low calcium intake and developing osteoporosis in this population.
Answer #1

Low Calcium

Yes

No

Total

Cases

50

100

150

Control

20

130

150

Total

70

230

300

Odds Ratio = (a/c)/(b/d) = (50/20)/(100/130) = 3.25

The odds are 3.25 times higher that a those who have low calcium will develop osteoporosis than those who do not low calcium.